OpenStudy (anonymous):

simplify the following expression express in terms of i square root of -64?

OpenStudy (anonymous):

8i

OpenStudy (saifoo.khan):

\[\Huge 8i\]

OpenStudy (anonymous):

how did u get that?

OpenStudy (saifoo.khan):

there is "i" because its -ve 64. because if -ve.

OpenStudy (anonymous):

\[i =\sqrt{-1}\] so: \[\sqrt{-64}=\sqrt{-1}*\sqrt{64}\] =i*8